OpenStudy (scarlettfarra2000):

Will someone please explain this to me? And help me I REALLY don't understand what I'm doing Evaluate each expression for the given values. Provide answers in simplified radical form and approximate irrational answers to the nearest hundredth.

OpenStudy (scarlettfarra2000):

\[15+2\sqrt{a+8} -\sqrt{3a}\] when a=4

OpenStudy (scarlettfarra2000):

\[2\sqrt{(x+y)-(x-y)^{2}}\] when x=8 and y=5

OpenStudy (scarlettfarra2000):

\[3\sqrt{(ab)^{2}}+(2a+b)^{3}\] when a =1 and b=4

OpenStudy (anonymous):

\[15+2\sqrt{a+8} -\sqrt{3a}\] for \(a=4\) start with \[15+2\sqrt{4+8} -\sqrt{3\times 4}\]

OpenStudy (anonymous):

there is more work to be done, but what do you get first?

OpenStudy (scarlettfarra2000):

\[17\sqrt{12}-\sqrt{12}\]

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

you can't add them like that, the first step is \[15+2\sqrt{12}-\sqrt{12}\]

OpenStudy (anonymous):

sort of like \[15+2x-x\] what do you get when you combine like terms there?

OpenStudy (scarlettfarra2000):

15+2x-x 2x

OpenStudy (anonymous):

where did the 15 go?

OpenStudy (scarlettfarra2000):

15+2x

OpenStudy (anonymous):

ok that is somewhat better but still not right what is \[2\times 5-5\]?

OpenStudy (scarlettfarra2000):

\[2\times5=10-5=5\]

OpenStudy (anonymous):

right and how about \[2\times 7-7\]?

OpenStudy (scarlettfarra2000):

\[2\times7=14-7=7\]

OpenStudy (anonymous):

ok good, so \[2\times 5-5=5\\2\times 7-7=7\\ 2\times x-x=?\]

OpenStudy (scarlettfarra2000):

5 and 7 would be the two x's?

OpenStudy (anonymous):

yes just trying to make sense out of \(2x-x\) for example \(x\) could be an apple two apples minus one apple is ?

OpenStudy (scarlettfarra2000):

if we had 2-1=1

OpenStudy (anonymous):

yes but is it not the number 2-1= 1 it is two apples minus one apple equals one apple \[2\spadesuit -\spadesuit=\spadesuit\] \[2x-x=?\]

OpenStudy (scarlettfarra2000):

the problem already gives us 2 \[2(2)-1=3?\] is that how we would set it up and solve?

OpenStudy (anonymous):

no, we are still at \[15+2\sqrt{12}-\sqrt{12}\]

OpenStudy (anonymous):

now the first question is, what is \(2\sqrt{12}-\sqrt{12}\)?

OpenStudy (scarlettfarra2000):

the factors of \[\sqrt{12}\] would be \[\sqrt{3} and \sqrt{4}\]and 4 can go smaller \[\sqrt{2}\]